# Agent Design Pattern Catalogue: A Collection of Architectural Patterns for Foundation Model based Agents

### v2

@@ A Collection of Architectural Patterns
Foundation model-enabled generative artificial intelligence facilitates the development and implementation of agents, which can leverage distinguished reasoning and language processing capabilities to takes a proactive, autonomous role to pursue users’ goals. Nevertheless, there is a lack of systematic knowledge to guide practitioners in designing the agents considering challenges of goal-seeking (including generating instrumental goals and plans), such as hallucinations inherent in foundation models, explainability of reasoning process, complex accountability, etc. To address this issue, we have performed a systematic literature review to understand the state-of-the-art foundation model-based agents and the broader ecosystem. In this paper, we present a pattern catalogue consisting of 17 architectural patterns with analyses of the context, forces, and trade-offs as the outcomes from the previous literature review. The proposed catalogue can provide holistic guidance for the effective use of patterns, and support the architecture design of foundation model-based agents by facilitating goal-seeking and plan generation.
@@ 1 Introduction
In this regard, we present a catalogue of patterns for the design of foundation model-based agents in this paper. Please note that “agent” can be referred to i) AI acting on behalf of another entity, or; ii) AI that can take active roles or produces effect to achieve users’ goals. The former circumstance requires thorough analysis on governance perspective, while hereby, we claim that in this study, we focus on the second concept of “agents” that are capable of goal-seeking and plan generation. In software engineering, an architectural pattern is a reusable solution to a problem that occurs commonly within a given context in software design. Our pattern catalogue includes 17 patterns that were identified based on the study conducted by Lu et al. [5]. The intended audience of collected patterns is software architects and developers who are interested in FM-based agent design and implementation. The contributions of this paper include:
@@ 2 Background & Related Work
The introduction of OpenAI’s ChatGPT [6] in November 2022 has gathered over 100 million users in two months upon its release, becoming the fastest-growing consumer internet app of all time [7]. This has also initiated the race among big tech arms in the development of FM and GenAI products. For instance, Google rolled out its own GenAI product, the Bard models  [^3] https://ai.google/static/documents/google-about-bard.pdf, then released the Gemini  [^4] https://gemini.google.com/ as the updated version. Anthropic has also emerged as one of the major LLM providers since the launch of their Claude models  [^5] https://www.anthropic.com/claude There are also many open-sourced FMs, such as Llama  [^6] https://llama.meta.com/ and Mistral  [^7] https://mistral.ai/.
@@
We noticed that there is a lack of a holistic view of architecture design, making it challenging to develop FM-based agents. Moreover, software architecture comprises software elements, relations among them, and properties of both [16]. With the increase in the incorporation of machine learning into software and systems, methods to identify the impact on the reliability of machine learning are essential to ensure the reliability of the software and systems in which these algorithms reside [17]. Hence, frameworks that only list the high-level components supporting their functionality usually lack system-level thinking, with no explicit identification of software components, relationships among them, and their properties [5]. Our work presents 17 design patterns regarding agent goal-seeking and plan generation. The pattern catalogue can provide a holistic guidance to practitioners on the trade-off analysis for the design of foundation model-based agents.
@@ 3 Methodology
Based on the reported findings, we delved into the analysis of identified patterns for building integrating FM-based agents. Through the SLR, we have identified a series of architectural design challenges in the development and implementation of systems with the integration of agents. We then further conducted extensive review on this topic, which includes grey literature, and real-world applications (through scrutinising official websites and documents available), for identifying the known uses. Combining the findings of both SLR and the additional review, in this paper, we report our findings on 17 extracted patterns.
@@
Agent evaluator Perform testing to assess the agent regarding diverse requirements and metrics, ensuring the functional suitability, adaptability with improved flexibility.
@@ 4 Pattern Catalogue for Foundation Model-based Agents
A generated plan should be reviewed to ensure its accuracy, usability, completeness, etc. Self-refection, cross-reflection, and human reflection can help the agent to collect feedback from different reflective entities, and refine the plan and reasoning steps accordingly. Afterwards, the agent can assign tasks to other narrow AI-based or non-AI systems, invoke external tools, and employ a set of agents for goal achievement by tool/agent registry. In particular, agents can work on the same task and finalise the results with voting-based, role-based, or debate-based cooperation. For instance, agents can act as different roles such as coordinator and worker. Multimodal guardrails can be applied to manage and control the inputs/outputs of foundation models. Please note that we omit the detailed architecture of agent-as-a-worker, and pattern application in several interactions for the simplicity of this diagram, for instance, the invocation of non-agent-AI/non-AI systems and external tools services can all apply tool/agent registry, and agent-as-a-worker can all have their respective datastore or knowledge base. Finally, we claim that developers can evaluate the performance of agents at both design-time and runtime via agent evaluator.
@@ 4.13 Role-based Cooperation
  • Li et al. [58] propose selective reflection-tuning in which the teacher agent’s reflection on data quality and student agent’s data selection are synergised.
@@ 4.16 Tool/Agent Registry
Solution: Fig. 16 depicts how can an agent search external agents and tools via a tool/agent registry. A user prompts goals to an agent, which then decomposes the goals into fine-grained tasks. The agent can query the tool/agent registry, which is the main entry point for collecting and categorising various tools and agents regarding a series of metrics (e.g., capability, price, context window). Based on the returned information, the agent can employ and assign the tasks to respective tools and agents. Please note that a registry can be implemented in different manners, for instance, a coordinator agent with specific knowledge based, blockchain-based smart contract, etc., and a registry can be extended into a marketplace for tool/agent service trading.
@@ 5 Discussion
Integration with extant patterns. Integrating different patterns can help structure and develop comprehensive and trustworthy agents. In particular, the proposed pattern catalogue can be applied together with responsible AI patterns [67] to ensure the agents behave in a responsible manner. For instance, bill of materials registry can record the outputs of components such as guardrails, prompt/response optimiser, etc., offering a complete supply chain of foundation model-based agent development. Whilst, black box can be implemented to collect the inputs and outputs of foundation model-based agents at runtime. If any abnormal behaviour is detected or an audit is needed, the recorded information can provide evidence for the accountability process. In addition, a set of voting mechanisms can be leveraged via blockchain smart contracts [68] for voting-based cooperation, to ensure transparent and secure interactions between agents.
@@ 6 Conclusion
Foundation model based agents are gaining increasing attention in various domains to intellectualise and automate the business process. However, practitioners are troubled by architectural challenges to design agents. Our previous work demonstrates a reference architecture to present an overview of agent design [5], while in this study, we scrutinise the forces, solutions, and trade-offs of 17 patterns. The pattern catalogue is provided as a holistic guidance for architects to better design and develop foundation model-based agents. In our future work, we will study how to apply the pattern catalogue with existing patterns to preserve the trustworthiness of agents, and further explore the architecture decisions that are related to foundation model-based agents.

### v3

@@ A Collection of Architectural Patterns
Foundation model-enabled generative artificial intelligence facilitates the development and implementation of agents, which can leverage distinguished reasoning and language processing capabilities to takes a proactive, autonomous role to pursue users’ goals. Nevertheless, there is a lack of systematic knowledge to guide practitioners in designing the agents considering challenges of goal-seeking (including generating instrumental goals and plans), such as hallucinations inherent in foundation models, explainability of reasoning process, complex accountability, etc. To address this issue, we have performed a systematic literature review to understand the state-of-the-art foundation model-based agents and the broader ecosystem. In this paper, we present a pattern catalogue consisting of 18 architectural patterns with analyses of the context, forces, and trade-offs as the outcomes from the previous literature review. The proposed catalogue can provide holistic guidance for the effective use of patterns, and support the architecture design of foundation model-based agents by facilitating goal-seeking and plan generation.
@@ 1 Introduction
In this regard, we present a catalogue of patterns for the design of foundation model-based agents in this paper. Please note that “agent” can be referred to i) AI acting on behalf of another entity, or; ii) AI that can take active roles or produces effect to achieve users’ goals. The former circumstance requires thorough analysis on governance perspective, while hereby, we claim that in this study, we focus on the second concept of “agents” that are capable of goal-seeking and plan generation. In software engineering, an architectural pattern is a reusable solution to a problem that occurs commonly within a given context in software design. Our pattern catalogue includes 18 patterns that were identified based on the study conducted by Lu et al. [5]. The intended audience of collected patterns is software architects and developers who are interested in FM-based agent design and implementation. The contributions of this paper include:
@@ 2 Background & Related Work
The introduction of OpenAI’s ChatGPT [6] in November 2022 has gathered over 100 million users in two months upon its release, becoming the fastest-growing consumer internet app of all time [7]. This has also initiated the race among big tech arms in the development of FM and GenAI products. For instance, Google rolled out its own GenAI product, the Bard models  [^3] https://ai.google/static/documents/google-about-bard.pdf, then released the Gemini  [^4] https://gemini.google.com/ as the updated version. Anthropic has also emerged as one of the major FM providers since the launch of their Claude models  [^5] https://www.anthropic.com/claude There are also many open-sourced FMs, such as Llama  [^6] https://llama.meta.com/ and Mistral  [^7] https://mistral.ai/. Schulhoff et al. [8] performed a systematic literature review and proposed a taxonomy of diverse prompting techniques for foundation models.
@@
We noticed that there is a lack of a holistic view of architecture design, making it challenging to develop FM-based agents. Moreover, software architecture comprises software elements, relations among them, and properties of both [17]. With the increase in the incorporation of machine learning into software and systems, methods to identify the impact on the reliability of machine learning are essential to ensure the reliability of the software and systems in which these algorithms reside [18]. Hence, frameworks that only list the high-level components supporting their functionality usually lack system-level thinking, with no explicit identification of software components, relationships among them, and their properties [5]. Our work presents 18 design patterns regarding agent goal-seeking and plan generation. The pattern catalogue can provide a holistic guidance to practitioners on the trade-off analysis for the design of foundation model-based agents.
@@ 3 Methodology
Based on the reported findings, we delved into the analysis of identified patterns for building integrating FM-based agents. Through the SLR, we have identified a series of architectural design challenges in the development and implementation of systems with the integration of agents. We then further conducted extensive review on this topic, which includes grey literature, and real-world applications (through scrutinising official websites and documents available), for identifying the known uses. Combining the findings of both SLR and the additional review, in this paper, we report our findings on 18 extracted patterns.
@@
Agent adapter Connect the agent and tools by learning new interfaces and converting incompatible interfaces into expected ones, ensuring interoperability and adaptability, and reduce development cost.
Agent evaluator Provide interface to connect the agent and external tools for task completion, ensuring the functional suitability, adaptability with improved flexibility.
@@ 4 Pattern Catalogue for Foundation Model-based Agents
A generated plan should be reviewed to ensure its accuracy, usability, completeness, etc. Self-refection, cross-reflection, and human reflection can help the agent to collect feedback from different reflective entities, and refine the plan and reasoning steps accordingly. Afterwards, the agent can assign tasks to other narrow AI-based or non-AI systems, invoke external tools, and employ a set of agents for goal achievement by tool/agent registry. In particular, agents can work on the same task and finalise the results with voting-based, role-based, or debate-based cooperation. For instance, agents can act as different roles such as coordinator and worker. Agent adapter keeps learning the interfaces of different tools, and convert them into FM-friendly environment. Multimodal guardrails can be applied to manage and control the inputs/outputs of foundation models. Please note that we omit the detailed architecture of agent-as-a-worker, and pattern application in several interactions for the simplicity of this diagram, for instance, the invocation of non-agent-AI/non-AI systems and external tools services can all apply tool/agent registry, and agent-as-a-worker can all have their respective datastore or knowledge base. Finally, we claim that developers can evaluate the performance of agents at both design-time and runtime via agent evaluator.
@@ 4.3 Prompt/Response Optimiser
  • Agent adapter. Prompt/response optimiser can improve users’ inputs, and the optimised prompts can be sent to other agents for goal achievement, while agent adapter focuses more on the utilisation of external tools.
@@ 4.13 Role-based Cooperation
  • Wang et al. [59] propose Mixture-of-Agents where proposer agents provide useful reference responses to aggregator agents, and the aggregator agents are composed in layers to synthesise and refine the responses.
@@ 4.16 Tool/Agent Registry
Solution: Fig. 16 depicts how can an agent search external agents and tools via a tool/agent registry. A user prompts goals to an agent, which then decomposes the goals into fine-grained tasks. The agent can query the tool/agent registry, which is the main entry point for collecting and categorising various tools and agents regarding a series of metrics (e.g., capability, price, context window). Based on the returned information, the agent can employ and assign the tasks to respective tools and agents. Please note that a registry can be implemented in different manners, for instance, a coordinator agent with specific knowledge base, blockchain-based smart contract, etc., and a registry can be extended into a marketplace for tool/agent service trading.
@@
  • Agent adapter. Tool/agent registry records the available external tools, while agent adapter can convert the interface of selected tools into agent-friendly format.
4.17 Agent Adapter
Summary: An agent adapter provides interface to connect the agent and external tools for task completion.
Context: An agent may leverage external tools to complete certain tasks for expanded capabilities.
Problem: The agent needs to deal with different interfaces of diverse tools, while certain interfaces might be incompatible or inefficient to interact for the agent. How can the agent assign tasks to external tools and process the results?
Forces:
  • Interoperability. Certain tasks require external tools to complete, and the tools may need agents to process particular information during intermediate steps.
  • Adaptability. Agents may employ new tools considering task complexity, tool capability, cost, etc.
  • Overhead. Manual development of compatible interfaces for agents and external tools can be intensive and inefficient.
Figure 17: Agent adapter.
Solution: Fig. 17 demonstrates a simplified graphical representation of agent adapter. Given user’s instructions, the agent generates a plan consisting of a set of tasks to achieve the user’s goals. In particular, the agent may employ diverse external tools to complete different tasks. However, tools have respective interfaces, which can be of different abstraction levels for the agent to deal with, or have specific format requirements, etc. Agent adapter can help invoke and manage these interfaces by converting the agent messages into required format or content, and vice versa. In particular, the adapter can retrieve tool manual or tutorial from datastore, to acquire available interfaces. It then transforms the agent outputs based on the interface requirements, and invokes the service [68]. Please note that fine-grained interface description can help agent to understand and hence improve the performance. The adapter also receives execution results from tools, which will be sent to the underlying foundation model for further analysis (e.g. task assignment to other tools, self-reflection for tool employment). For instance, the adapter can translate tasks into system messages when interacting with local file system, or capture and operate graphical user interface when playing a video game.
Consequences:
Benefits:
  • Interoperability. Agent adapter facilitates the interoperation between an agent and external tools.
  • Adaptability. Agents can employ new tools via agent adapter, which can acquire and convert the tool API via corresponding manual or tutorial.
  • Reduced development cost. Agent adapter enables autonomous conversion of interfaces, there is no need to develop compatible interfaces for different tools, hence the development cost is reduced.
Drawbacks:
  • Maintenance overhead. i) Agent adapter itself requires proper maintenance and evaluation to ensure the correctness of outputs. ii) Agent adapter may need additional memory or external data store to record the historical tool interfaces.
Known uses:
  • AutoGen  [^22] https://microsoft.github.io/autogen/docs/tutorial/tool-use. Users can register different tools in the agent, specifying the usage description. Registered tools will be leveraged by the agent during a conversation with user.
  • Apple Intelligence  [^23] https://www.apple.com/apple-intelligence/. Apple Intelligence can support writing, image generation, schedule management across different products and applications. For instance, it can capture the entities in users’ photo library and create emoji.
  • Semantic Kernel  [^24] https://learn.microsoft.com/en-us/semantic-kernel/agents/plugins/?tabs=Csharp. Semantic Kernel can orchestrate agents and plugins to extend agents’ skills. Plugins need to provide semantic description (e.g. input, output, side effects) for agents to understand.
  • Yang et al. [69] devise SWE-agent that can provide agent-computer interfaces, enabling foundation model-based agents to process code commands and resolve software engineering tasks.
Related patterns:
  • Prompt/response optimiser. Prompt/response optimiser can improve users’ inputs, and the optimised prompts can be sent to other agents for goal achievement, while agent adapter focuses more on the utilisation of external tools.
  • Tool/agent registry. Tool/agent registry records the available external tools, while agent adapter can convert the interface of selected tools into agent-friendly format.
@@ 5 Discussion
Integration with extant patterns. Integrating different patterns can help structure and develop comprehensive and trustworthy agents. In particular, the proposed pattern catalogue can be applied together with responsible AI patterns [70] to ensure the agents behave in a responsible manner. For instance, bill of materials registry can record the procurement of components such as guardrails, prompt/response optimiser, etc., offering a complete supply chain of foundation model-based agent development. Whilst, black box can be implemented to collect the inputs and outputs of foundation model-based agents at runtime. If any abnormal behaviour is detected or an audit is needed, the recorded information can provide evidence for the accountability process. In addition, a set of voting mechanisms can be leveraged via blockchain smart contracts [71] for voting-based cooperation, to ensure transparent and secure interactions between agents.
@@ 6 Conclusion
Foundation model based agents are gaining increasing attention in various domains to intellectualise and automate the business process. However, practitioners are troubled by architectural challenges to design agents. Our previous work demonstrates a reference architecture to present an overview of agent design [5], while in this study, we scrutinise the forces, solutions, and trade-offs of 18 patterns. The pattern catalogue is provided as a holistic guidance for architects to better design and develop foundation model-based agents. In our future work, we will study how to apply the pattern catalogue with existing patterns to preserve the trustworthiness of agents, and further explore the architecture decisions that are related to foundation model-based agents.